OpenStudy (anonymous):

can some one give me a good example and explanation on how to write an equation of a line with a slope

OpenStudy (phi):

This video and the ones following it might answer your question http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/algebra--slope

OpenStudy (anonymous):

When noting the equaltion of a line you generally have 2 options, the first being in point slope form. \[(y-y0)=m(x-x0)\] This form is often used and left as is when noting the equation for a tangent line, with m being the slope and x0 and y0 being replaced by their respective values. This can be simplified into the second, and most recognizable, option for the equation of a line, the Slope intercept form: \[f(x)=mx+b\] This form presents the equation as a function of x, with m being a constant for the slope and b being a constant for the value of the function when x=0 (the value of the function when it's graph crosses the vertical axis. Both notations are equally acceptable!